Poulter Corporation will pay a dividend of $3.25 per share next year. The company pledges to increase its dividend by 5.1 percent per year, indefinitely. If you require a return of 11 percent on your investment, how much will you pay for the company’s stock today?
Answer #1

Next year's dividend = D1 = $3.25

Dividend growth rate = g = 5.1% (indefinitely)

Required rate of return = r = 11%

We can calculate the price of the company's stock today using Gordon's Growth Model

According to Gordon's growth model, price of a stock P is give by:

P = D1/(r - g)

P = 3.25/(11% - 5.1%) = $55.0847457627119

Answer -> current price of the stock = 55.0847457627119